# The VLLM Safety Paradox: Dual Ease in Jailbreak Attack and Defense

### v1

@@ 6 Vision-free Evaluator
Table 5: ASR of four LLMs w and w.o an explicit system protection prompt defense on the VLSafe and FigStep datasets. For VLSafe, we omit the protection prompt as it lacks specific scenario contexts defined (see Table 1).
@@
We propose two variants of vision-free evaluators: i) the Naive method, which directly takes the original textual instruction only as input; ii) the Scenario method, which includes a scenario-aware system protection prompt (e.g., potentiality to expose to illegal activity) before the given instruction. Four recent LLMs serve as candidate evaluators: Mistral-7B [24], QWen2.5-14B [52], Llama3.1-8B, and Llama3.1-70B [14].
6.1 Results on Jailbreak Datasets
The overall results for the four evaluators are presented in Table 5 and Table 6 (values on the left side of symbol $\rightarrow$). From these tables, we draw the following two observations:
  • The vision-free evaluators already achieve highly competitive results on jailbreak datasets. For instance, as shown in Table 5, Llama3.1-8B reaches an ASR of 0.7 on VLSafe, while Llama3.1-70B with the Scenario variant achieves an ASR of 2.8 on FigStep. Exceptions observed on VLGuard and MM-SafetyBench (Table 6) stem from instructions requiring joint image-text understanding.
  • The Scenario approach consistently outperforms its Naive counterpart by a significant performance margin in most cases. This finding suggests that informing LLMs explicitly about the potential for harmful scenarios enhances their confidence in identifying jailbreak instances.
Caption Re-check. We note that queries from the other two jailbreak datasets, VLGuard [71] and MM-SafetyBench [33], demand a joint understanding of both image and instruction. To address the limitations of LLMs lacking access to visual information, we propose using QWen2-VL-14B [56] to generate captions for the provided images, enabling LLMs to utilize these captions as contextual background.
Table 6 presents the results before and after the caption integration step, separated by $\rightarrow$. It can be observed that i) most models exhibit a decreasing trend in ASR, indicating that captions, particularly those containing OCR-embedded information, can reveal harmful content recognized by LLMs. ii) One exception is the QWen2.5 model, which shows a notable increase in ASR. We delve into the generated responses of this model and find that QWen2.5 often declines to answer harmful queries, though without using the standard keywords typically defined in  [71]. This finding further highlights the limitations of current evaluation metrics, underscoring the need for a more robust and comprehensive metric in this domain.
@@ 6.1 Results on Jailbreak Datasets
6.2 Helpfulness and Safety Trade-off
Building on the above results, we thereby propose an LLM Pipeline approach to balancing model response safety and helpfulness. This approach follows a two-step pipeline: first, an instruction is evaluated by an LLM (e.g., Llama3.1). If it passes the safety check, it is then input to a VLLM for response generation. We evaluate this method’s performance on two LLaVA-1.5 models, comparing it against two defense-aware strategies [^3] We use LLaVA-1.5 models because  [71] provides only fine-tuned checkpoints for these models.. Additionally, we leverage the Safe-Safe and Safe-Unsafe categories from VLGuard for this experiment [71]. Specifically, Safe-Safe is evaluated using the winning rate metric, estimated by GPT-4o [2], while Safe-Unsafe is evaluated based on the ASR.
The results are presented in Fig. 9. From this figure, we observe that: i) while the vanilla LLaVA-1.5 models perform best in the Safe-Safe category, they make substantial compromises in defense effectiveness; ii) the defense-aware PostHoc approach experiences a significant drop in performance within the Safe-Safe category. It is worth noting that the PostHoc approach [71] has already been fine-tuned on the tested dataset. In contrast, our proposed LLM-Pipeline method achieves the best trade-off between model safety and helpfulness. Moreover, we delve into cases that are labeled as safe by Llama3.1 from the Safe-Unsafe category. Fig. 10 presents two such examples, and we find these two pose minimal actual risk. This finding highlights the dataset limitation and underscores the need to collect more representative instances or establish a clearer, consensus-based definition of safety in multi-modalities.
@@ 7 Conclusion and Discussion
Summary. This work presents a worrisome safety paradox within existing VLLMs. We conduct an in-depth study of both sides of jailbreak attacks and defenses, tentatively revealing the underlying rationales for these two, particularly the issue of over-prudence in current defense mechanisms. In addition, we propose repurposing existing LLM guardrails to function as a vision-free evaluator as a potential alternative solution.
Future directions:
  • Modern applications of (V)LLMs are no longer limited to standalone models. Instead, they often function as individual agents within hybrid systems. Compared to explicit malicious content, scenarios involving hybrid information structures present more complex attack dimensions, such as imperceptible toxic triggers, prompt injection [47] and long-context jailbreaking [^4] https://www.anthropic.com/research/many-shot-jailbreaking.. Consequently, developing benchmarks tailored to these scenarios can better unveil the vulnerability of modern (V)LLMs.
  • On the defense side, Reinforcement Learning deserves further research attention, as even simple rule-based rewards have shown significant promise [37]. Second, system-level strategies, such as prioritizing system instructions to mitigate prompt injection, contribute another promising direction. Moreover, distilling safety alignment capabilities from LLMs appears to be a more efficient strategy than developing defense methods for VLLMs from scratch.

### v2

@@ 5.1 Detector Results on Jailbreak Datasets
Table 5: ASR of four LLMs on the VLSafe and FigStep datasets. Scenario refers to the inclusion of an additional system protection prompt before the given instruction. For VLSafe, we omit the protection prompt as it lacks specific scenario contexts (see Table 1).
@@
Table 6 presents the results before and after the caption integration step, separated by $\rightarrow$. It can be observed that i) most models exhibit a decreasing trend in ASR, indicating that captions, particularly those containing OCR-embedded information, can reveal harmful content recognized by LLMs. ii) One exception is the QWen2.5 model, which shows a notable increase in ASR. We delve into the generated responses of this model and find that QWen2.5 often declines to answer harmful queries, though without using the standard keywords typically defined in  [75].
@@
5.2 Detect-then-Respond: LLM-Pipeline
Building on the above results, we thereby design an LLM Pipeline approach to balancing model response safety and helpfulness. This approach follows a two-step pipeline: 1) An instruction is evaluated by an LLM detector (i.e., Llama3.1). 2) If it passes the safety check, it is then input to a VLLM for response generation; otherwise, the query will be rejected. We evaluate this method’s performance on two LLaVA-1.5 models, comparing it against two defense-aware strategies [^3] We use LLaVA-1.5 models because  [75] provides only fine-tuned checkpoints for these models.. Additionally, we utilize the Safe-Safe and Safe-Unsafe categories from VLGuard [75], which are intended to be answered and rejected, respectively Specifically, Safe-Safe is evaluated using the winning rate metric (helpfulness), estimated by GPT-4o [2], while Safe-Unsafe is evaluated based on ASR (harmlessness).
The results are presented in Fig. 9. From this figure, we observe that: i) while the vanilla LLaVA-1.5 models perform best in the Safe-Safe category, they make substantial compromises in defense effectiveness; ii) the defense-aware PostHoc approach experiences a significant drop in performance within the Safe-Safe category. It is worth noting that the PostHoc approach [75] has already been fine-tuned on the tested dataset. In contrast, our proposed LLM-Pipeline method achieves a better trade-off between model harmlessness and helpfulness. Moreover, we delve into cases that are labeled as safe by Llama3.1 from the Safe-Unsafe category. Fig. 10 presents two such examples, and we find these two pose minimal actual risk. This finding highlights the dataset limitation and underscores the need to collect more representative instances or establish a clearer, consensus-based definition of safety in multi-modalities.
@@ 5.2 Detect-then-Respond: LLM-Pipeline
6 Related Work
6.1 VLLM Attack
Existing jailbreak attacks on VLLMs can be broadly categorized into two groups: adversarial perturbation and prompt injection [32, 35]. The former involves optimizing an adversarial image [40], either from random noise or a benign image, to elicit harmful responses [6, 53, 62, 50, 72]. The objective of this attack is to generate outputs that include a predefined list of toxic words. For instance, [21, 47] show that a single visual adversarial input can universally jailbreak an aligned VLLM. In contrast to these methods that operate within a constrained perturbation budget, prompt injection techniques deliberately manipulate image or instruction data without such limitations [10, 39, 65, 75]. The dominant techniques in this category focus on embedding high-risk content into images through typography or generative methods like stable diffusion [49]. For example, FigStep [19] utilizes textual prompts to induce MLLMs into completing sentences in an image that inadvertently result in malicious outputs step-by-step. MM-SafetyBench [36] generates harmful images spanning 13 commonly encountered scenarios. SASP [65] aims to hijack the system prompt by using GPT-4 [42] as a red teaming tool against itself, searching for potential jailbreak prompts.
6.2 VLLM Defense
6.3 LLM Attack and Defense
LLM jailbreak attack methods can be roughly classified into white-box and black-box attacks based on the transparency of the victim models [76, 66, 7]. White-box attack strategies include efforts to search for jailbreak prompts by leveraging model gradients [27, 74, 76] or predicted logits of output tokens [71, 23]. Additionally, some methods involve fine-tuning the target LLMs with adversarial examples to induce harmful behaviors [29, 48, 70]. In contrast, prompt manipulation constitutes the primary method employed in the more challenging black-box attacks [69, 15, 67]. To defend against such jailbreak attacks, various approaches have been proposed, including safeguarding system prompts [52, 77], implementing supervised fine-tuning [4, 5, 14], and developing RLHF techniques [3, 43, 17, 58].
@@ 7 Conclusion and Discussion
Summary. This work presents a worrisome safety paradox within existing VLLMs. We conduct an in-depth study of both sides of jailbreak attacks and defense, tentatively revealing the underlying rationales for these two, particularly the issue of over-prudence in current defense mechanisms. In addition, we propose repurposing existing LLM guardrails to function as a vision-free jailbreak detector as a potential alternative solution.
It is important to note that the LLM-Pipeline method is not intended to serve as a better jailbreak defense baseline, as there is minimal to no room for improvement. Instead, we leverage this approach to underscore the uncertainty in this field: rather than focusing efforts on designing a sophisticated VLLM defense mechanism, the advanced built-in LLM guardrails already help yield favorable results. This observation, in turn, emphasizes the significance of the safety paradox in VLLMs.
Future directions. Building on the insights from this work, we outline the following three directions, i.e., attack, defense, evaluation, that deserve more attention in the future:
  • Collection of comprehensive attack dataset. Modern applications of (V)LLMs are no longer limited to standalone models. Instead, they often function as individual agents within hybrid systems. Compared to explicit malicious content, scenarios involving hybrid information structures present more complex attack dimensions, such as imperceptible toxic triggers, prompt injection [51] and long-context jailbreaking [^4] https://www.anthropic.com/research/many-shot-jailbreaking.. Consequently, developing benchmarks tailored to these scenarios can better unveil the vulnerability of modern (V)LLMs.
  • Development of robust defense method. On the defense side, Reinforcement Learning deserves further research attention, as even simple rule-based rewards have shown significant promise [41]. Second, system-level strategies, such as prioritizing system instructions to mitigate prompt injection, contribute another promising direction. Moreover, distilling safety alignment capabilities from LLMs appears to be a more efficient strategy than developing defense methods for VLLMs from scratch.